# Simulating and investigating various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model

### v1

@@ IV.3 The effects of dissipations on evolution
In the first row of Fig. 7, when the initial state is $|0\rangle|0\rangle|1\rangle$, result is found that the larger dissipations intensities $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$, the larger the probability of forming stable hydrogen bond. However, in the second row, when initial state is $|1\rangle|1\rangle|0\rangle$, the opposite result is found. In the first row from panel (a) to panel (d), it is easy to find that the solid dividing line representing the probability of $0.9$ is shifted from top to bottom until it disappears. At the same time, the area of region III of probability with [0.5, 0.9)0.50.9[0.5,\ 0.9)[ 0.5 , 0.9 ) decreases to 00 and the area of region IV of probability with [0.9, 1]0.91[0.9,\ 1][ 0.9 , 1 ] increases to $1$. This means the intensity $\gamma_{phn}^{out}$ promotes the hydrogen bond formation when initial state is $|0\rangle|0\rangle|1\rangle$. Similarly, in the second row from panel (e) to panel (h), the solid dividing line representing the probability of $0.1$ is shifted from right to left. At the same time, the area of region I of probability with [0, 0.1)00.1[0,\ 0.1)[ 0 , 0.1 ) increases and the area of region II of probability with [0.1, 0.5]0.10.5[0.1,\ 0.5][ 0.1 , 0.5 ] decreases. This means intensity $\gamma_{phn}^{out}$ hinders the hydrogen bond formation when initial state is $|1\rangle|1\rangle|0\rangle$. In summary, in the case of $|0\rangle|0\rangle|1\rangle$, the all dissipations intensities have the positive effect on the formation of stable hydrogen bond; in the case of $|1\rangle|1\rangle|0\rangle$, the dissipations intensities have the negative effect on the formation of stable hydrogen bond. Similar to the Fig. 6, when the initial state is $|0\rangle|0\rangle|1\rangle$, the probabilities on the heat maps are all on the range of [0.5, 1]0.51[0.5,\ 1][ 0.5 , 1 ]; When the initial state is $|1\rangle|1\rangle|0\rangle$, the probabilities on the heat maps are all on the range of [0, 0.5)00.5[0,\ 0.5)[ 0 , 0.5 ).
@@ IV.4 The effects of inflows on evolution
Temperature is also an important factor, it will cause inflows of phonons. According to Eq. (5), $\mu_{bond},\ \mu_{isol}$ and $\mu_{phn}$ are described as follows
@@
where $\omega_{isol}=\omega_{bond}=\omega_{phn}$, $\mu_{isol}<1,\ \mu_{bond}<1,\ \mu_{phn}<1$. When $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$, $\mu_{bond},\ \mu_{isol}$ and $\mu_{phn}$ are positively correlated with inflows intensities $\gamma_{isol}^{in}$, $\gamma_{bond}^{in}$ and $\gamma_{phn}^{in}$, respectively; they are also positively correlated with temperature $T_{isol}$, $T_{bond}$ and $T_{phn}$, respectively. For the convenience of research, $\mu_{isol}$, $\mu_{bond}$ and $\mu_{phn}$ can be used to study the effects of inflows or temperatures on evolution.
According to the result in Fig. 8, no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$, the larger $\mu_{isol}$ and the smaller $\mu_{bond}$, the larger the probability of forming stable hydrogen bond. This means $\mu_{isol}$ promotes the hydrogen bond formation, but $\mu_{bond}$ hinders the hydrogen bond formation. In the first row from panel (a) to panel (d), it is easy to find that all three solid dividing lines representing the probability of $0.1$, $0.5$ and $0.9$ respectively are shifted from left to right. At the same time, the area of region I of probability with [0, 0.1)00.1[0,\ 0.1)[ 0 , 0.1 ) and the area of region II of probability with [0.1, 0.5)0.10.5[0.1,\ 0.5)[ 0.1 , 0.5 ) both increases, the area of region III of probability with [0.5, 0.9]0.50.9[0.5,\ 0.9][ 0.5 , 0.9 ] and the area of region IV of probability with [0.9, 1]0.91[0.9,\ 1][ 0.9 , 1 ] both decreases. This means the $\mu_{phn}$ hinders the hydrogen bond formation no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$. In summary, the $\mu_{isol}$ has the positive effect on the formation of stable hydrogen bond; however, the $\mu_{bond}$ and $\mu_{phn}$ have the negative effect on it. Different to the Figs. 6 and 7, the probabilities on the heat maps are all on the range of [0, 1]01[0,\ 1][ 0 , 1 ].
@@ IV.5 The effects of external impulses on evolution
Figure 9: (online color) Effect of external impulses of phonons on the hydrogen bond formation. $g_{dist}=g_{prot}=g$, $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{ph}^{out}=\gamma$, $\mu_{isol}=\mu_{bond}=0$, $\mu_{phn}=0.01$. The number of phonons varies in different panels: 10 in panel (a), 20 in (b), 30 in (c).
@@
In Fig. 10, the stability of hydrogen bonds increases with the increase of the number of phonons. When the number of phonons increases, the probability of a stable hydrogen bond increases, while the probability of a broken hydrogen bond decreases, and both tend to a stable value.
Figure 11: (online color) Effect of the initial phonon number and phonon inflow on the hydrogen bond formation. $g_{dist}=g_{prot}=g$. The initial state is $|0\rangle|0\rangle|N_{phn}\rangle$. In the first row, $\gamma_{isol}^{out}=\gamma_{phn}^{out}$ is fixed, while $\gamma_{bond}^{out}$ is varied in the four panels: $\gamma_{bond}^{out}$ = 0.1 in panel (a), $\gamma_{bond}^{out}$ = 0.4 in panel (b), $\gamma_{bond}^{out}$ = 0.7 in panel (c), $\gamma_{bond}^{out}$ = 1 in panel (d). In the second row, $\gamma_{bond}^{out}=\gamma_{phn}^{out}$ is fixed, while $\gamma_{isol}^{out}$ is varied in the four panels: $\gamma_{isol}^{out}$ = 0.1 in panel (e), $\gamma_{isol}^{out}$ = 0.4 in panel (f), $\gamma_{isol}^{out}$ = 0.7 in panel (g), $\gamma_{isol}^{out}$ = 1 in panel (h).
In Fig.11 the effect of external impulse and phononic temperature on the probability of stable hydrogen bond formation is obtained. From the horizontal axis, as the external impulse (the number of phonons in the initial state) increases, the hydrogen bond stability decreases. From the vertical axis, the same conclusion applies to the phononic temperature $\mu_{ph}$. Their increase will make the micro-oscillations of proton more intense, thereby reducing the stability of hydrogen bonds.
From the four panels (a),(b),(c),(d) in the first row, as $\gamma_{bond}^{out}$ increases, the temperature of the heat map increases, indicating that it is positively correlated with the stability of hydrogen bonds. In addition, by comparing the columns of these four pictures, the increase in $\gamma_{bond}^{out}$ also reduces the effect of $\mu_{phn}$ on the probability of stable hydrogen bond formation, making the temperature of the heat map gradually approach inversely proportional to the number of phonons. From the four panels in the second row, as $\gamma_{isol}^{out}$ increases, the temperature of the heat map decreases, indicating that it is negatively correlated with the stability of hydrogen bonds. By comparing panel (e) with panels (f), (g), and (h), when $\gamma_{isol}^{out}$ is small enough, the temperature of the heat map hardly changes with the increase in the number of phonons, indicating that the probability of stable hydrogen bonds is high.
@@ V Concluding discussion and future work
  • In Sec. IV.2, the effect of interactions on evolution is studied. When the initial state is $|0\rangle|0\rangle|1\rangle$, the larger the interaction strengths, the harder to form a hydrogen bond. On the contrary, when the initial state is $|1\rangle|1\rangle|0\rangle$, the larger the interaction strengths, the easier to form a hydrogen bond.
  • In Sec. IV.3, the effect of dissipations on the hydrogen bond formation is studied. When the initial state is $|0\rangle|0\rangle|1\rangle$, the probability of stable hydrogen bond formation is negatively correlated with $\gamma_{isol}^{out}$ and positively correlated with $\gamma_{bond}^{out}$; when the initial state is $|1\rangle|1\rangle|0\rangle$, the probability of stable hydrogen bond formation is negatively correlated with both $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$.
  • In Sec. IV.4, the effect of inflows on the hydrogen bond formation is studied. Regardless of whether the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$, the probability of stable hydrogen bond formation is positively correlated with $\mu_{isol}$ and negatively correlated with $\mu_{bond}$.
  • In Sec. IV.5, the effect of external impulses on the hydrogen bond formation is obtained. The more free phonons flow in from the outside, the lower the probability of stable hydrogen bonds forming, and as the number of phonons increases, the probability approaches a critical value.
The results show that the evolution and hydrogen bond formation can be controlled by selectively choosing system parameters. Although we only studied the dynamic aspects of simple hydrogen bond model, the results we found will be used as a basis to extend the research to more complex chemical and biological model in the future.
@@ Appendix A Quantum master equation for the target model
The QME has the following form
@@
where $\mathcal{L}=L^{in}+L^{out}$. And
@@ Appendix B Estimation of physical quantities
The average energy of hydrogen bonds in water is $21\ kJ/mol$. To convert this energy to the energy of a single hydrogen bond, estimate it by using Avogadro’s constant ($N_{A}$), the number of particles per mole of substance, which is approximately ($6.022\times 10^{23}$). Through calculation, it can be estimated that the energy of a single hydrogen bond
@@
Substituting into Eq. (21) and calculating, the characteristic time scale is got as follows
$\tau=\frac{\hbar}{E}=\frac{\hbar\times N_{A}}{E_{H-bond}}=\frac{1.0545718\times 10^{-34}\ [J\cdot s]\times 6.022\times 10^{23}\text{[molecules/mol}]}{21\times 10^{3}\ \text{[J/mol]}}\approx 3.175316\times 10^{-15}\ [s]$ (29)

### v2

@@ IV.3 The effects of dissipations on evolution
In the first row of Fig. 7, when the initial state is $|0\rangle|0\rangle|1\rangle$, result is found that the larger dissipations intensities $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$, the larger the probability of forming stable hydrogen bond. However, in the second row, when initial state is $|1\rangle|1\rangle|0\rangle$, the opposite result is found. In the first row from panel (a) to panel (d), it is easy to find that the solid dividing line representing the probability of $0.9$ is shifted from top to bottom until it disappears. At the same time, the area of region III of probability with [0.5, 0.9)0.50.9[0.5,\ 0.9)[ 0.5 , 0.9 ) decreases to 00 and the area of region IV of probability with [0.9, 1]0.91[0.9,\ 1][ 0.9 , 1 ] increases to $1$. This means the intensity $\gamma_{phn}^{out}$ promotes the hydrogen bond formation when initial state is $|0\rangle|0\rangle|1\rangle$. Similarly, in the second row from panel (e) to panel (h), the solid dividing line representing the probability of $0.1$ is shifted from right to left. At the same time, the area of region I of probability with [0, 0.1)00.1[0,\ 0.1)[ 0 , 0.1 ) increases and the area of region II of probability with [0.1, 0.5]0.10.5[0.1,\ 0.5][ 0.1 , 0.5 ] decreases. This means intensity $\gamma_{phn}^{out}$ hinders the hydrogen bond formation when initial state is $|1\rangle|1\rangle|0\rangle$. In summary, in the case of $|0\rangle|0\rangle|1\rangle$, the all dissipations intensities have the positive effect on the formation of stable hydrogen bond; in the case of $|1\rangle|1\rangle|0\rangle$, the dissipations intensities have the negative effect on the formation of stable hydrogen bond. Similar to the Fig. 6, when the initial state is $|0\rangle|0\rangle|1\rangle$, the probabilities on the heat maps are all on the range of [0.5, 1]0.51[0.5,\ 1][ 0.5 , 1 ]; When the initial state is $|1\rangle|1\rangle|0\rangle$, the probabilities on the heat maps are all on the range of [0, 0.5)00.5[0,\ 0.5)[ 0 , 0.5 ). This shows that dissipations intensities have a slight effect on evolution, too.
@@ IV.4 The effects of inflows on evolution
Temperature is also an important factor. According to Eq. (7), $\mu_{bond},\ \mu_{isol}$ and $\mu_{phn}$ are described as follows
@@
where $\omega_{isol}=\omega_{bond}=\omega_{phn}$, $\mu_{isol}<1,\ \mu_{bond}<1,\ \mu_{phn}<1$. When $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$, $\mu_{bond},\ \mu_{isol}$ and $\mu_{phn}$ are positively correlated with inflows intensities $\gamma_{isol}^{in}$, $\gamma_{bond}^{in}$ and $\gamma_{phn}^{in}$, respectively; they are also positively correlated with temperature $T_{isol}$, $T_{bond}$ and $T_{phn}$, respectively. For the convenience of research, $\mu_{isol}$, $\mu_{bond}$ and $\mu_{phn}$ can be used to study the effects of inflows or temperatures on evolution. The values of $\mu_{isol}$ and $\mu_{bond}$ range from 00 to $0.9$. The values of $\mu_{phn}$ from the first column to the fourth column correspond to 0, 0.3, 0.600.30.60,\ 0.3,\ 0.60 , 0.3 , 0.6 and $0.9$, respectively. Under different initial states, the dissipative evolution will also eventually reach probabilistic stability after long-term evolution.
Figure 9: (online color) Effect of external impulses of phonons on the hydrogen bond formation. $g_{dist}=g_{prot}=g$, $\gamma_{isol}^{out}=\gamma_{bond}^{out}=\gamma_{phn}^{out}=\gamma$, $\mu_{isol}=\mu_{bond}=0$, $\mu_{phn}=0.01$. The number of phonons varies in different panels: 10 in panel (a), 20 in (b), 30 in (c).
According to the Fig. 8, result is found that no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$, the larger $\mu_{isol}$ and the smaller $\mu_{bond}$, the larger the probability of stable hydrogen bond formation. This indicates that whereas $\mu_{bond}$ inhibits the creation of hydrogen bonds, $\mu_{isol}$ encourages them. In the first row from panel (a) to panel (d), it is easy to find that all three solid dividing lines representing the probability of $0.1$, $0.5$ and $0.9$ respectively are shifted from left to right. At the same time, the area of region I of probability with [0, 0.1)00.1[0,\ 0.1)[ 0 , 0.1 ) and the area of region II of probability with [0.1, 0.5)0.10.5[0.1,\ 0.5)[ 0.1 , 0.5 ) both increases, the area of region III of probability with [0.5, 0.9]0.50.9[0.5,\ 0.9][ 0.5 , 0.9 ] and the area of region IV of probability with [0.9, 1]0.91[0.9,\ 1][ 0.9 , 1 ] both decreases. This means the $\mu_{phn}$ hinders the hydrogen bond formation no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$. In summary, the $\mu_{isol}$ has the positive effect on the formation of stable hydrogen bond; however, the $\mu_{bond}$ and $\mu_{phn}$ have the negative effect on it. Different to the Figs. 6 and 7, the probabilities on the heat maps are all on the range of [0, 1]01[0,\ 1][ 0 , 1 ]. This shows that inflows intensities (temperature) have a significant effect on evolution than that of other factors (interactions and dissipations).
@@ IV.5 The effects of external impulses on evolution
Now consideration is given to the effect of both the initial phonon number and the phonon inflow intensity on the evolution. In Fig. 11, the effect of external impulse and phononic temperature (replaced by $\mu_{phn}$) on the probability of stable hydrogen bond formation is obtained. From the horizontal axis, as the external impulse (the number of phonons in the initial state) increases, the hydrogen bond stability decreases. From the vertical axis, the same conclusion applies to the phononic inflow intensity $\mu_{phn}$. The increase of temperature will make the micro-oscillations of proton more intense, thereby reducing the stability of hydrogen bonds. From the four panels of the first row in Fig. 11, as $\gamma_{bond}^{out}$ increases, the probability of the points in the heat map increases, indicating that $\gamma_{bond}^{out}$ promotes the formation of stable hydrogen bond. In addition, the larger $\gamma_{bond}^{out}$ is, the less obvious the negative effect of $\mu_{phn}$ on stable hydrogen bond formation is. From the four panels of the second row in Fig. 11, as $\gamma_{isol}^{out}$ increases, the probability of the points in the heat map decreases, indicating that $\gamma_{bond}^{out}$ hinders the formation of stable hydrogen bond. In panel (e), $\gamma_{isol}^{out}$ is small enough, the the probability of the points in the heat map hardly changes with the increase of the number of phonons; in panels (f)$\sim$(h), the values of $\gamma_{isol}^{out}$ are large, the probability of the points in each heat map changes with the increase of the number of phonons.
@@ V Concluding discussion and future work
  • In Sec. IV.2, the effect of interactions on evolution is studied. When the initial state is $|0\rangle|0\rangle|1\rangle$, the larger the interaction strengths, the harder to form a hydrogen bond, indicating that the interactions hinders the formation of stable hydrogen bond. On the contrary, when the initial state is $|1\rangle|1\rangle|0\rangle$, the larger the interaction strengths, the easier to form a hydrogen bond, indicating that the interactions promotes the formation of stable hydrogen bond. The addition of weak inflows promotes the formation of hydrogen bond, regardless of whether the initial state is $|1\rangle|1\rangle|0\rangle$ or $|0\rangle|0\rangle|1\rangle$.
  • In Sec. IV.3, the effect of dissipations on the hydrogen bond formation is studied. When the initial state is $|0\rangle|0\rangle|1\rangle$, the probability of stable hydrogen bond formation is negatively correlated with $\gamma_{isol}^{out}$ and positively correlated with $\gamma_{bond}^{out}$; however, when the initial state is $|1\rangle|1\rangle|0\rangle$, the probability of stable hydrogen bond formation is negatively correlated with both $\gamma_{isol}^{out}$ and $\gamma_{bond}^{out}$. $\gamma_{phn}^{out}$ promotes the hydrogen bond formation when the initial state is $|0\rangle|0\rangle|1\rangle$, and hinders it when the initial state is $|1\rangle|1\rangle|0\rangle$.
  • In Sec. IV.4, the effect of inflows on the hydrogen bond formation is studied. Regardless of whether the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$, the probability of stable hydrogen bond formation is positively correlated with $\mu_{isol}$ and negatively correlated with $\mu_{bond}$. $\mu_{phn}$ hinders the formation of hydrogen bond no matter the initial state is $|0\rangle|0\rangle|1\rangle$ or $|1\rangle|1\rangle|0\rangle$.
  • In Sec. IV.5, the effect of external impulses on the hydrogen bond formation is obtained. It is evident that the more free phonons flow in system from the outside, the lower the probability of stable hydrogen bonds formation, and as the number of phonons increases, the probability approaches a limit. The same conclusion as in the previous subsection is obtained: $\mu_{phn}$ hinders the formation of hydrogen bond in the case of large number of phonons.
[…]
Based on the above conclusions, our work demonstrates that the evolution and hydrogen bond formation can be controlled by selectively choosing system parameters. In particular, they are greatly affected by temperature. Although we only studied the dynamic aspects of simple hydrogen bond model, the results we found will be used as a basis to extend the research to more complex chemical and biological model in the future.
@@ Appendix A Quantum master equation for the target model
According to the previous sections, the QME has the following form
@@
where $H$ is Hamiltonian, $\rho$ is density matrix, $\mathcal{L}=L^{in}+L^{out}$. Then
@@ Appendix B Estimation of physical quantities
The average energy of hydrogen bonds in water is $E_{bond}=21[\rm{kJ/mol}]$. To convert this energy to the energy of a single hydrogen bond, estimate it by using Avogadro’s constant $N_{A}$ — the number of particles per mole of substance, which is approximately ($6.022\times 10^{23}$). Through calculation, it can be estimated that the energy of a single hydrogen bond
@@
Replacing $E$ in Eq. (23) with $E_{H-bond}$, the characteristic time scale is got as follows
$\tau=\frac{\hbar}{E_{H-bond}}=\frac{\hbar\times N_{A}}{E_{bond}}=\frac{1.0545718\times 10^{-34}\ [J\cdot s]\times 6.022\times 10^{23}\text{[molecules/mol}]}{21\times 10^{3}\ \text{[J/mol]}}\approx 3.175316\times 10^{-15}\ [\rm{s}]$ (31)